2. Let us assume that the population X has the mean μ and the variance σ2 and the population Y h 2σ. If X and Y are independent, express the following quantities by and ơ as the mean u and the variance (2.2) V[X-Y] (2.3) V[2X+3Y] (2.4) VIX-3Y-5]
Answer #1

Question 2

Here as we know that

E[aX + bY] = a E[X] + b E[Y]

V[aX + bY] = a2V[x] + b2V[Y]

V[aX - bY] = a2V[x] + b2V[Y]

(a) E[X -2] = E[X] - 2 = \mu -2

(b) V[X - Y] = V[X] + V[Y] = \sigma^2 + 2 * \sigma^2 = 3\sigma^2

(c) V[2X + 3Y] = 22V[X] + 32 V[Y] = 4\sigma^2 + 9 * 2\sigma^2 = 22 \sigma^2

(d) V[X - 3Y - 5] = V[X] + 32 V[Y] = \sigma^2 + 9 * 2 \sigma^2 = 19 \sigma^2